OpenStudy (anonymous):

Sketch a graph of the following function on paper. f(x) = 2 sin(1/3x − π/3) Find the zero(s) of f on the interval [0 + θ, T + θ), where T is the period and θ is the phase shift. (Enter your answers as a comma-separated list. If an answer does not exist, enter DNE.)

OpenStudy (xapproachesinfinity):

if you to sketch that graph you need to find the period first \[w=\frac{2\pi}{|B|}\] if the function is \[f(x)=A\sin(Bx-C)\]

OpenStudy (xapproachesinfinity):

so in your case the period is \[w=\frac{2\pi}{1/3}=6\pi\]

OpenStudy (xapproachesinfinity):

so every 6 pi this function will repeat it self we also need to know that sin is odd function (symmetrical with respect to the origin)

OpenStudy (xapproachesinfinity):

of course you need to find the zeros then look for some points after than you can draw the graph

OpenStudy (xapproachesinfinity):

do you know how to find the zeros

OpenStudy (xapproachesinfinity):

?

OpenStudy (anonymous):

No, I don't understand how to find zeroes @xapproachesinfinity

OpenStudy (xapproachesinfinity):

you have to set f(x)=0 meaning you need to solve \[2\sin(\frac{x}{3}-\frac{\pi}{3})=0 \Longrightarrow \sin(\frac{x}{3}-\frac{\pi}{3})=0\]